# Balanced and Elastic End-to-end Training of Dynamic LLMs

### v1

@@ 3.3.1. Overhead of DynMo and Frequency of Dynamism
As a general rule, we apply rebalancing via DynMo whenever the model or control flow changes. Given its low overhead, DynMo can be used as frequently as needed, depending on application requirements. For example, in gradual pruning, dynamism (that is, model changes requiring rebalancing) typically occurs every few thousand iterations. In contrast, for MoEs and MoDs, rebalancing is needed every iteration due to unpredictable imbalances caused by routing decisions taken during the forward pass at each FFN. In these cases, we perform rebalancing during backpropagation, coupling layer migration with the pipeline parallelism scheme by moving layers as gradients propagate from the last to the first layer.
3.4. Re-packing Model to Fewer Workers
@@
7:         if mem_usage[src] + mem_usage[dst] $<$ MAX_MEM
8:    && sum(active_workers) $>$ target_num_workers then
@@ 3.4.2. Releasing Unused GPUs after Re-packing
To ensure the GPUs are released in a practical manner, we use the NCCL communicator. While the NCCL communicator can not be resized, NCCL supports splitting the communicator via ncclCommSplit() into multiple sub-partitions, or even creating a single communicator with fewer ranks (Nvidia, 2024). After re-packing, active GPUs are assigned to a new sub-communicator, and idle GPUs to a separate one. Since idle GPUs are excluded from the active communicator, concurrent communicators can safely proceed without the risk of deadlock.
Alternatively, re-packing can be coordinated with checkpointing. Unlike rebalancing, which may occur every few iterations, re-packing is infrequent—typically triggered after thousands of iterations, when the model structure has changed significantly. By combining re-packing with a checkpoint restart, implementation is simplified, since a new NCCL communicator is already created during the restart. Moreover, because the model is reloaded and reshared among the workers during checkpoint recovery, there is no additional overhead for resharding the model to a new set of workers.
@@
Lastly, even though DynMo’s speedup mostly comes from dynamic load balancing, re-packing provides 4–11% additional performance benefit from the reduced communication.
@@ 4. Implementation
The gather and scatter operations in global pruning were implemented by employing NCCL Peer-to-Peer (P2P) send-receive operations instead of collectives since the sizes of the objects to be sent (local_topk) and received (indices_to_keep) in Algorithm 1 from each rank are different and other ranks do not have this size information to participate in the collective call.
@@ 4.2.1. Mixture of Experts
We build on Mixtral 8x7b weights from Hugging Face (Mistral, 2024) and LLaMA-MoE-3.5B (Team, 2024) in continual training by monitoring the imbalance between layers with respect to number of assigned tokens. We used both the auxiliary load balance scheme adopted in Mixtral 8x7b (Jiang et al., 2024) and the S-BASE load balancer (Lewis et al., 2021b). Since the imbalance is dependent on the point the routing decision is taken, i.e., in the forward pass at each FFN, we rebalance in the back propagation phase where we attach the movement of layers to the pipeline parallelism.
@@ 4.2.2. Parameter Pruning
Among the six example cases, gradual pruning posed a unique challenge, as it required code changes to integrate with DynMo. Unstructured pruning demands a sparse storage format for efficient training and transfer. A common choice is the compressed sparse row (CSR) format, which necessitates replacing dense matrix multiplications (DMM) with sparse equivalents (SpMM). Since PyTorch does not support computing the derivative of SpMM operations for backpropagation on a CSR tensor, we evaluated GPU-compatible CSR-based SpMM implementations, namely cuSPARSE by Nvidia and Sputnik (Gale et al., 2020). Sputnik’s SpMM consistently outperformed cuSPARSE across all tested sparsity levels. This is largely due to Sputnik’s kernels being tailored for deep learning workloads, unlike cuSPARSE, which is optimized for HPC workloads with extreme sparsity (often ¿99%). Notably, Sputnik begins to outperform cuBLAS around 75% sparsity. Thus, for sparse operations, we implemented PyTorch bindings for the CUDA kernels of Sputnik [^3] The Sputnik bindings are made available at the following link: https://anonymous.4open.science/r/Torch-Sputnik-E926/README.md..
@@ 4.2.6. Mixture of Depths
We build on the MoDs work by Raposo et al. (Raposo et al., 2024) by including in our GPT models we use in testing the small auxiliary MLP predictor that predicts whether that token will be among the top-k or not in the sequence. Similar to the case of MoEs, since the imbalance is dependent on the point the router takes the decision, i.e. in the forward pass, we rebalance in the back propagation phase where we attach the movement of layers to the pipeline parallelism scheme. Since the routing happens around the entire block, i.e., the routing applies to both to both forward MLPs and multi-head attention, we treat the skipped layers to be shadow layers when redistributing the layers on workers.
@@ 5. Evaluation
Experiments were primarily conducted on a supercomputer where each compute node consists of 2× AMD EPYC 9654 96-core/2.4GHz CPUs and 4× NVIDIA H100 SXM5 (80GB) GPUs. GPUs communicate with CPUs over PCIe Gen5 x16, and with each other via NVSwitch (NVLink34 ×6). Nodes are interconnected via 4× 200Gbps InfiniBand NDR200 links. We used CUDA 12.6, OpenMPI 4.0.7, and PyTorch 2.3.1 with the NCCL 2.17.1 backend.
Models were trained on the Wikipedia dataset (Foundation, 2023). For Mixtral 8×7B and LLaMA-MoE-3.5B, we perform continual training. All models use a sequence length of 2048, hidden size of 1024, and 32 attention heads. Unless otherwise specified, training runs for 10,000 iterations with micro-batch size 2 and batch size 64. In multi-node experiments, as we scale the number of GPUs, we proportionally increase the batch size to maintain four micro-batches per GPU, following the guidance in (Huang et al., 2019) for optimal pipeline utilization.
@@
Figure 4. Re-packing the layers of GPT models into fewer GPUs as the model gets smaller due to dynamism. X-axis: Number of GPUs in pipeline parallelism using 90 nodes and up to 8 GPUs per node. Left Y-axis: throughput/number of GPUs shown with yellow bars. Right Y-axis: throughput (tokens/sec) shown with blue squares. Below: we show the average number of GPUs needed throughout the gradual pruning training at which we dynamically re-pack (total 10,000 iterations). Right: load balancing overhead for example cases. The overhead reported is broken down to three things: profiling (in XX), DynMo load balancing algorithm (in XX), and migration of layers between GPUs (in XX).
@@ 5.1. End-to-end Training Throughput and Speedup
We observed that the use of layer execution time for dynamic load balancing, such as diffusion or partitioning, consistently outperforms parameter count-based implementations across all scales. In every scale, execution time-based dynamic balancers surpass the baseline static balancers. As seen in the figure, most of the speedup reported is attributed to balancing the load by redistributing the, and not due to the reduced communication when we re-pack: the speedup gain from re-packing is between 4$\sim$11of the entire speedup gain. In other words, even if we do not re-pack at, the speedup gain remains almost the same. Hence we treat re-packing as just a way to be efficient by using less GPUs, and not as a way to speed up imbalanced dynamic training.
@@ 5.3. Re-packing Models to Fewer GPUs
5.4. Additional Results and Ablation
Additional results and ablation are included in the supplementary material allowed for accepted papers: effect of change of network bandwidth on load balancing, single node vertical scaling, multi-node weak scaling, and dynamic minibatch/microbatch size.
@@ 6.1. Load Balancing Model-Parallel DNNs
Load balancing via graph partitioning. Graph partitioning-based load balancing schemes identify atomic operations in the model and represent them as nodes in a directed acyclic graph (DAG), where edges capture data dependencies. Tanaka et al.(Tanaka et al., 2021) partition the DAG in three phases: identifying atomic operations, grouping them into blocks based on computation time, and combining these blocks into final partitions using a dynamic programming-based algorithm. Qararyah et al.(Qararyah et al., 2021) form disjoint clusters by identifying critical paths and mapping them to devices using an algorithm that accounts for both communication minimization and temporal load balancing. Both approaches rely on profiling before training and perform partitioning only once.
@@ 7. Conclusion
The primary aim and impact, we argue, is that research on dynamic models will not deliver practical impact unless there is a platform from which those models can be made efficient. DynMo serves as that platform, offering a load-balancing system and a re-packing solution specifically designed for dynamic models whose worker loads vary during training. This leads to improved efficiency and faster end-to-end training times. Empirical results on LLMs, demonstrated across six representative dynamic model cases, show that DynMo significantly improves training throughput compared to existing approaches. Moreover, it enables elastic GPU utilization, resulting in additional cost savings. As dynamic models become increasingly prevalent, dynamic load distribution will be critical for scalable and efficient training.

### v2

@@ 3.3.1. Overhead of DynMo and Frequency of Dynamism
As a general rule, we apply rebalancing via DynMo whenever the model or control flow changes. Given its low overhead, DynMo can be used as frequently as needed, depending on application requirements. For example, in gradual pruning, dynamism (that is, model changes requiring rebalancing) typically occurs every few thousand iterations. In contrast, for MoEs and MoDs, rebalancing is needed every iteration due to unpredictable imbalances caused by routing decisions taken during the forward pass at each FFN. In these cases, we perform rebalancing during backpropagation, coupling layer migration with the pipeline parallelism scheme by moving layers while the gradients calculation take place, from the last to the first layer.
3.4. Re-packing the Model to Fewer Workers
@@
7:   if mem_usage[src] + mem_usage[dst] << MAX_MEM
8:    && sum(active_workers) >> target_num_workers then
@@ 3.4.2. Releasing Unused GPUs after Re-packing
To ensure the GPUs are released in a practical manner, we use the NCCL communicator. While the NCCL communicator can not be resized, NCCL supports splitting the communicator via ncclCommSplit() into multiple sub-partitions, or even creating a single communicator with fewer ranks (Nvidia, 2024). After re-packing, active GPUs are assigned to a new sub-communicator, while idle GPUs are assigned to a different one. Since idle GPUs are excluded from the active communicator, concurrent communicators can safely proceed without the risk of deadlock.
Alternatively, re-packing can be coordinated with checkpointing. Unlike rebalancing, which may occur every few iterations, re-packing is infrequent—typically triggered after thousands of iterations, when the model structure has changed significantly. By combining re-packing with a checkpoint restart, the implementation is simplified since a new NCCL communicator is already created during the restart. Moreover, because the model is reloaded and reshared among the workers during checkpoint recovery, there is no additional overhead for resharding the model to a new set of workers.
@@
Lastly, even though DynMo’s speedup mostly comes from dynamic load balancing, re-packing provides 4–11% additional performance benefit from the reduced number of number of ranks and volume of communication.
@@ 4. Implementation
The gather and scatter operations in global pruning were implemented by employing NCCL Peer-to-Peer (P2P) send-receive operations instead of collectives since the sizes of the objects to be sent (local_topk) and received (indices_to_keep) in Algorithm 1 from each rank are different and other ranks do not have this size information to participate in the collective call. It is worth mentioning that while an alltoallv collective could be used as an alternative, using it however would add unnecessary global synchronization and may cause contention. Thus we opt for P2P communication.
@@ 4.2.1. Mixture of Experts
We build on Mixtral 8x7b weights from Hugging Face (Mistral, 2024) and LLaMA-MoE-3.5B (Team, 2024) in continual training by monitoring the imbalance between layers with respect to number of assigned tokens. We used both the auxiliary load balance scheme adopted in Mixtral 8x7b (Jiang et al., 2024) and the S-BASE load balancer (Lewis et al., 2021b). Since the imbalance is dependent on the point the routing decision is taken, i.e., in the forward pass at each FFN, we rebalance in the back propagation phase where we attach the movement of layers to the pipeline parallelism communication flow.
@@ 4.2.2. Parameter Pruning
Among the six example cases, gradual pruning posed a unique challenge, as it required code changes to integrate with DynMo. Unstructured pruning demands a sparse storage format for efficient training and transfer. A common choice is the compressed sparse row (CSR) format, which necessitates replacing dense matrix multiplications (DMM) with sparse equivalents (SpMM). Since PyTorch does not support computing the derivative of SpMM operations for backpropagation on a CSR tensor, we evaluated GPU-compatible CSR-based SpMM implementations, namely cuSPARSE by Nvidia and Sputnik (Gale et al., 2020). Sputnik’s SpMM consistently outperformed cuSPARSE across all tested sparsity levels. This is largely due to Sputnik’s kernels being tailored for deep learning workloads, unlike cuSPARSE, which is optimized for HPC workloads with extreme sparsity (often ¿99%). Notably, Sputnik begins to outperform cuBLAS around 75% sparsity. Thus, for sparse operations, we implemented PyTorch bindings for the CUDA kernels of Sputnik.
@@ 4.2.6. Mixture of Depths
We build on the MoDs work by Raposo et al. (Raposo et al., 2024) by including in our GPT models we use in testing the small auxiliary MLP predictor that predicts whether that token will be among the top-k or not in the sequence. Similar to the case of MoEs, since the imbalance is dependent on the point the router takes the decision, i.e. in the forward pass, we rebalance in the back propagation phase where we attach the movement of layers to the pipeline parallelism scheme. Since the routing happens around the entire block, i.e., the routing applies to both forward MLPs and multi-head attention, we treat the skipped layers to be shadow layers when redistributing the layers on workers.
@@ 5. Evaluation
Experiments were primarily conducted on a supercomputer where each compute node consists of 2x AMD EPYC 9654 96-core/2.4GHz CPUs and 4x NVIDIA H100 SXM5 (80GB) GPUs. GPUs communicate with CPUs over PCIe Gen5 x16, and with each other via NVSwitch (NVLink34 x6). Nodes are interconnected via 4x 200Gbps InfiniBand NDR200 links. We used CUDA 12.6, OpenMPI 4.0.7, and PyTorch 2.3.1 with the NCCL 2.17.1 backend.
Models were trained on the Wikipedia dataset (Foundation, 2023). For Mixtral 8x7B and LLaMA-MoE-3.5B, we perform continual training. All models use a sequence length of 2048, hidden size of 1024, and 32 attention heads. Unless otherwise specified, training runs for 10,000 iterations with micro-batch size 2 and batch size 64. In multi-node experiments, as we scale the number of GPUs, we proportionally increase the batch size to maintain four micro-batches per GPU, following the guidance in (Huang et al., 2019) for optimal pipeline utilization.
@@
Figure 4. Re-packing the layers of GPT models into fewer GPUs as the model gets smaller due to dynamism. X-axis: Number of GPUs in pipeline parallelism using 90 nodes and up to 8 GPUs per node. Left Y-axis: throughput/number of GPUs shown with yellow bars. Right Y-axis: throughput (tokens/sec) shown with blue squares. Below: we show the average number of GPUs needed throughout the gradual pruning training at which we dynamically re-pack (total 10,000 iterations). Right: load balancing overhead for example cases. The overhead reported is broken down to three things: profiling (in $X$), DynMo load balancing algorithm (in $X$), and migration of layers between GPUs (in $X$).
@@ 5.1. End-to-end Training Throughput and Speedup
[…]
@@ 6.1. Load Balancing Model-Parallel DNNs
Load balancing via graph partitioning. Graph partitioning-based load balancing schemes identify atomic operations in the model and represent them as nodes in a directed acyclic graph (DAG), where edges capture data dependencies. Tanaka et al.(Tanaka et al., 2021) partition the DAG in three phases: identifying atomic operations, grouping them into blocks based on computation time, and combining these blocks into final partitions using DP. Qararyah et al.(Qararyah et al., 2021) form disjoint clusters by identifying critical paths and mapping them to devices using an algorithm that accounts for both communication minimization and temporal load balancing. Both approaches rely on profiling before training and partition only once.
@@ 7. Conclusion
Research on dynamic models will not deliver practical impact unless there is a platform from which those models can be made efficient. DynMo serves as that platform, offering a load-balancing system and a re-packing solution specifically designed for dynamic models whose worker loads vary during training. This leads to improved efficiency and faster end-to-end training times. Empirical results on LLMs, demonstrated across six representative dynamic model cases, show that DynMo significantly improves training throughput compared to existing approaches. Moreover, it enables elastic GPU utilization, resulting in additional cost savings. As dynamic models become increasingly prevalent, dynamic load distribution will be critical for scalable and efficient training.
Acknowledgements.
Authors from Koç University has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 949587).